OpenStudy (anonymous):

The area of a rectangular counter top is w2 – 3w – 10. Which of the following could be the length of the counter top? A. w – 10 B. w – 3 C. w + 2 D. w + 5

OpenStudy (anonymous):

i know its asking alot but just put the answer please

OpenStudy (anonymous):

(w-5)(w+2) if you factorise so since AREA=length*width then length=w+2 width=w-5

OpenStudy (anonymous):

makes that C